# MS-0001-1922.65

### v1

@@ 3.2 Performance Evaluation of the Machine Learning Models
Among the variables corresponding to the characteristics of applicant devices, medical specialties, country codes, and product codes are highly predictive of the recall risk. The results are consistent with prior literature indicating some heterogeneous effects of these indicators on the risk of recall. Furthermore, our results show that Life Sustaining/Supporting is of significant importance. This variable quantifies the risk level associated with an applicant device. Riskier applicant devices are generally under stringent market scrutiny and are more likely to be recalled. Finally, with respect to the similarity between an applicant device and its predicate devices, we observe that neither Num. of Unmatched Specialties nor Num. of Unmatched Product Codes are among the top 15 most informative variables, suggesting that taking them into account will not be that useful.
@@
Note: US = Unites States; DE = Germany ; CN = China ; KR = Korea; SU = General & Plastic Surgery
@@ 4 Design of a Data-Driven Human-Algorithm Decision Support Tool
We develop an advanced clearance policy by harnessing the power of an ML model. However, in cases where the predicted risk for an applicant device may not be informative enough, our policy has the capability to offer supplementary information for assessment by human experts without presenting a direct recommendation. The evaluation of such intricate cases is delegated to the designated FDA committee, leveraging their expertise to arrive at more informed judgments. The supplementary information are provided as risk labels, which can be generated using the predicted risks by our ML model.
@@ 4.1 A Data-Driven Advanced Clearance Policy
Our policy has two main components: (1) an ML predictor to estimate recall risk of a 510(k) device based on the information available to the FDA upon submission by the manufacturer, and (2) an optimization approach that determines whether an applicant device can be quickly accepted/rejected or if it should be deferred and more elaborately evaluated by an FDA assigned committee. Additionally, a classifier can be used to categorize deferred 510(k) devices based on their risk level, thereby providing more supporting information for the assigned committee to guide their decision-making. By combining these main core elements, our policy streamlines the clearance process, ensuring that devices are assessed with greater efficiency and accuracy while there are enough resources for in-depth evaluation of deferred devices which require more attention.
In our policy, the evaluation occurs through two main phases, as depicted in Figure 6. In the initial phase, the ML model discussed in §3.1 assesses the recall risk of an applicant device based on the available information provided during the submission. If the predicted risk by the ML model is lower than an optimized low threshold ($\ell$), the policy recommends accepting the device. If, however, the predicted risk exceeds the high threshold ($h$), the device is recommended to be rejected. The policy optimizes the specific values of the low and high thresholds to minimize the rates of acceptance of unsafe devices and rejection of safe devices, while still aligning with the preferences of decision-makers. That is, it ensures that (a) rates of rejection of unsafe devices and acceptance of safe devices are at least greater than values specified as desired, and (b) the workload imposed to the FDA for more elaborate evaluation of deferred devices does not exceed a preferred level. For devices that are deferred to a FDA committee for more in-depth evaluation, the policy assigns a risk label, ranging from very low risk to very high risk (the number of risk labels illustrated in Figure 6 are for illustrative purposes and can be refined as needed). These risk labels provide additional information to assist the committees in making better judgments.
@@
As mentioned, the first core element of our policy is predicting the recall risk for an applicant device given the information available upon application submission. In Section §3.1, we discussed how we have developed and trained a well-performing ML model using our dataset. Given a vector of inputs containing information on an applicant device and its corresponding predicate devices, the model generates a probability of recall. Let $f:X\rightarrow(0,1)$ be the functional representation of the model, which maps the vector of information on an applicant device and its predicates denoted by $X$ to a probability value in $(0,1)$. The second core element of our policy contains developing a constraint optimization model to identify the low and high thresholds. Let $\delta^{+}=f(X|Y=1)$ be the recall risk estimated for an applicant device which has been recalled according to data ($Y=1$). Similarly, let $\delta^{-}=f(X|Y=0)$ be the recall risk estimated for an applicant device with no recall in data ($Y=0$). Using this notation, we next describe the logic behind the optimization model that forms the second core element of our proposed policy.
A well-designed and effective policy is expected to yield low rates of acceptance of unsafe devices $(\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell)])$ and rejection of safe devices $(\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h)])$. However, a low value for those measures may yield low rates of rejection of unsafe devices $(\mathbb{E}[\mathbbm{1}(\delta^{+}\geq h)])$ and acceptance of safe devices $(\mathbb{E}[\mathbbm{1}(\delta^{-}\leq\ell)])$. Also, a low workload $(\mathbb{E}[\mathbbm{1}(\ell<f(X)<h)])$ for the FDA’s committees is desired so only a low ratio of applicant devices should be deferred and judged by them. Due to trade-offs between these measures, identifying the values of low and high risk thresholds ($\ell$ and $h$, respectively) is challenging. We develop a non-linear optimization model to solve this challenging problem:
$\min_{\ell,h}$ $\;\;\lambda\;\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell)]+(1-\lambda)\;\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h)]$ (1)
s.t. $\;\;\mathbb{E}[\mathbbm{1}(\delta^{+}\geq h)]\;\geq\;\xi^{ru}$ (2)
$\;\;\mathbb{E}[\mathbbm{1}(\delta^{-}\leq\ell)]\;\geq\;\xi^{as}$ (3)
$\;\;\mathbb{E}[\mathbbm{1}(\ell<f(X)<h)]\;\leq\;\rho$ (4)
@@ 4.2 Structural Properties
The optimization problem (1)-(5) is challenging to solve due to its non-convexity. In this section, we derive important structural properties of the optimization problem, allowing us to find a closed-form solution in many instances and develop an efficient heuristic to find near-optimal results in other instances.
@@
$\min_{\ell,h}$ $\;\;\lambda\;{\color[rgb]{0,0,0}\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell)]}+(1-\lambda)\;{\color[rgb]{0,0,0}\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h)]}$ (6)
s.t. $\;\;{\color[rgb]{0,0,0}\mathbb{E}[\mathbbm{1}(\delta^{+}\geq h)]\;\geq\;\xi^{ru}}$ (7)
$\;\;{\color[rgb]{0,0,0}\mathbb{E}[\mathbbm{1}(\delta^{-}\leq\ell)]\;\geq\;\xi^{as}}$ (8)
@@
To solve the Relaxed Problem, we introduce an auxiliary problem for which we can derive a closed-form solution. We then demonstrate how this auxiliary problem is in essence equivalent to the relaxed problem. The auxiliary problem is:
@@
$\max_{\ell,h}$ $\;-\theta\;\ell+(1-\theta)\;h$ (10)
[…]
where $\theta\in(0,1)$.
The following lemma demonstrates the equivalence between a linear program and the Auxiliary Problem, which is non-linear in general. This equivalence, in turn, allows deriving a closed-form solution for the Auxiliary Problem.
Lemma 4.1 (LP Equivalence)
The Auxiliary Problem is equivalent to the following linear program:
@@
where $h(\xi^{ru})\;=\;\sup\left\{h\in[0,1]:\;\mathbb{P}(\delta^{+}\geq h)\;\geq\;\xi^{ru}\right\}$ denotes the highest value of $h$ for which the true negative rate is greater than or equal to $\xi^{ru}$, and $\ell(\xi^{as})\;=\;\inf\left\{\ell\in[0,1]:\;\mathbb{P}(\delta^{-}\leq\ell)\;\geq\;\xi^{as}\right\}$ denotes the smallest value of $\ell$ for which the true positive rate is greater than or equal to $\xi^{as}$.
Using Lemma 4.1, we next derive a closed-form solution for the Auxiliary Problem.
@@
    if $h(\xi^{ru})\;>\;\ell(\xi^{as})$, then the two-thresholds optimal solution of the Auxiliary Problem is optimal in the Relaxed Problem,
@@
4.3 Lagrangian Relaxation
Given the structural properties derived in the prior section, the optimization problem proposed in §4.1 is still challenging to solve due to the workload constraint, which couples the computations of the low and high thresholds. In this section, we present a Lagrangian relaxation method that effectively relaxes the constraints related to workload and thresholds feasibility. To handle violations of the workload constraints and infeasible thresholds, we incorporate Lagrangian penalty terms into the formulation. Lagrangian relaxations have been extensively explored in the literature (see e.g., Brown and Smith 2020, Brown and Zhang 2022, and Liu et al. 2022). A key innovation in the Lagrangian relaxation we consider is the capability to simplify the optimization problem significantly by decomposing it into two straightforward sub-problems.
We start off by defining the primal problem which is equivalent to the main optimization proposed in §4.1:
Primal Problem
$\min_{\ell,h}$ $\;\;\lambda\;\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell)]+(1-\lambda)\;\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h)]$
s.t. $\;\;h\;\leq\;h(\xi^{ru})$
$\;\;\ell\;\geq\;\ell(\xi^{as})$
$\;\;\phi(h)-\phi(\ell)\;\leq\;p$
[…]
where $\phi(y)=\mathbb{P}(f(X)\leq y)$ is the empirical cumulative distribution function (CDF) of $f(X)$. We note that the workload constraint can also be computed as a function of the four indicators introduced in §4.1. Nevertheless, we continue using the empirical CDF for notation simplicity.
Next, we define the Lagrangian problem by dropping the FDA’s workload constraint and the feasibility requirement on the low and high thresholds, and by penalizing their violation in the objective function with Lagrangian multipliers $\gamma_{1}$ and $\gamma_{2}$, respectively.
Lagrangian Primal Problem
For a fixed $\gamma_{1}\geq 0$ and $\gamma_{2}\geq 0$:
$\min_{\ell,h}$ $\;\;\mathcal{L}(\ell,h,\gamma_{1},\gamma_{2})\;=\;\lambda\;\mathbb{E}[\mathbbm{1}(\delta^{+}\leq\ell)]+(1-\lambda)\;\mathbb{E}[\mathbbm{1}(\delta^{-}\geq h)]+\gamma_{1}\;(\phi(h)-\phi(\ell)-p)+\gamma_{2}\;(\ell-h)$
s.t. $\;\;h\;\leq\;h(\xi^{ru})$
$\;\;\ell\;\geq\;\ell(\xi^{as})$
[…]
The following result provides some useful properties of the Lagrangian Primal Problem.
Proposition 4.4
For any non-negative Lagrange multipliers $\gamma_{1}$ and $\gamma_{2}$:

### v2

@@ 3.1 Performance Evaluation of the Machine Learning Models
Among the variables corresponding to the characteristics of applicant devices, medical specialties, country codes, and product codes are highly predictive of the recall risk. The results are consistent with prior literature indicating some heterogeneous effects of these indicators on the risk of recall. Furthermore, our results show that Life Sustaining/Supporting is of significant importance. This variable quantifies the risk level associated with an applicant device. Riskier applicant devices are generally under stringent market scrutiny and are more likely to be recalled. Finally, with respect to the similarity between an applicant device and its predicate devices, we observe that neither Prop. of Unmatched Specialties nor Prop. of Unmatched Product Codes are among the top 15 most informative variables, suggesting that taking them into account will not be that useful.
To validate our key findings regarding important predictors of recall risk, we examine the logistic regression model with Lasso, which achieves performance comparable to our selected model. Specifically, we evaluate the $p$-values from a refitted logistic regression model that includes only the variables selected by Log Reg with Lasso (see Appendix 10). The results confirm that our findings are robust, and the interpretability provided by the SHAP analysis is highly consistent with the statistical inference offered by linear models. Regarding the variables Predicate Median Age, Predicate Oldest, and Variance of Recalls, although they contribute to improved predictive performance in the selected model, the results suggest that their association with recall risk should be interpreted with caution.
@@
Note: US = United States; DE = Germany; CN = China; RA = Radiology; KR = Korea; CH = Clinical Chemistry
@@ 4 Design of a Data-Driven Human-Algorithm Decision Support Tool
We develop an advanced clearance policy by harnessing the power of an ML model. However, in cases where the predicted risk for an applicant device may not be informative enough, our policy has the capability to offer supplementary information for assessment by human experts without presenting a direct recommendation. The evaluation of such intricate cases is delegated to the designated FDA committee, leveraging their expertise to arrive at more informed judgments.
@@ 4.1 A Data-Driven Advanced Clearance Policy
Our policy has two main components: (1) an ML predictor to estimate the recall risk of a 510(k) device based on the information available to the FDA upon submission by the manufacturer, and (2) an optimization approach that determines whether an applicant device can be accepted/rejected or if it should be deferred and more elaborately evaluated by an FDA-assigned committee. For deferred devices, the estimated risk can be provided to the FDA-assigned committee as additional information to assist the committee in making better judgments. Although risk labels (e.g., low, moderate, high) could be assigned to each deferred device, doing so may introduce additional complexity and subjectivity in defining the label thresholds. For devices recommended for acceptance or rejection by our policy, we envision a minimal level of human review to ensure that the final decision remains under FDA’s oversight. This modest review could involve a brief verification process that requires only a small fraction of the effort typically needed for a full FDA evaluation. By combining these main core elements, our policy streamlines the clearance process, ensuring that devices are assessed with greater efficiency and accuracy while there are enough resources for in-depth evaluation of deferred devices that require more attention.
In our policy, the evaluation occurs through two main phases, as depicted in Figure 8. In the initial phase, the ML model discussed in Section 3 assesses the recall risk of an applicant device based on the available information provided during the submission. If the predicted risk by the ML model is lower than an optimized low threshold ($\ell$), the policy recommends accepting the device. If, however, the predicted risk exceeds the high threshold ($h$), the device is recommended to be rejected. The policy optimizes the values of the low and high thresholds to minimize the rates of acceptance of unsafe devices and rejection of safe devices, while still aligning with the preferences of decision-makers. That is, it ensures that (a) rates of rejection of unsafe devices and acceptance of safe devices are at least greater than values specified as desired and (b) the workload imposed on the FDA for more elaborate evaluation of deferred devices does not exceed a preferred level.
@@
As mentioned, the first core element of our policy is predicting the recall risk for an applicant device given the information available upon application submission. In Section 3.1, we discussed how we have developed and trained a well-performing ML model using our dataset. Given a vector of inputs containing information on an applicant device and its corresponding predicate devices, the model generates a probability of recall. Let $f:X\rightarrow(0,1)$ be the functional representation of the model, which maps the vector of information on an applicant device and its predicates denoted by $X$ to a probability value in $(0,1)$. This function approximates the conditional probability $\mathbb{P}(Y=1|X)$. The second core element of our policy contains developing a constraint optimization model to identify the low and high thresholds. Let $\delta^{+}=f(X)|Y=1$ be the recall risk estimated for an applicant device which has been recalled according to data ($Y=1$). Similarly, let $\delta^{-}=f(X)|Y=0$ be the recall risk estimated for an applicant device with no recall in data ($Y=0$). Note that both $\delta^{+}$ and $\delta^{-}$ are random variables as they are functions of $X$. Using this notation, we next describe the logic behind the optimization model that forms the second core element of our proposed policy.
A well-designed and effective policy is expected to yield low rates of acceptance of unsafe devices $(\mathbb{P}(\delta^{+}\leq\ell))$ and rejection of safe devices $(\mathbb{P}(\delta^{-}\geq h))$. However, a low value for those measures may yield low rates of rejection of unsafe devices $(\mathbb{P}(\delta^{+}\geq h))$ and acceptance of safe devices $(\mathbb{P}(\delta^{-}\leq\ell))$. Also, a low workload $(\mathbb{P}(\ell<f(X)<h))$ for the FDA’s committees is desired so only a low ratio of applicant devices should be deferred and judged by them. Due to trade-offs between these measures, identifying the values of low and high risk thresholds ($\ell$ and $h$, respectively) is challenging. We develop a non-linear optimization model to solve this challenging problem:
$\min_{\ell,h}$ $\;\;\lambda\;\mathbb{P}(\delta^{+}\leq\ell)+(1-\lambda)\;\mathbb{P}(\delta^{-}\geq h)$ (1)
s.t. $\;\;\mathbb{P}(\delta^{+}\geq h)\;\geq\;\xi^{ru}$ (2)
$\;\;\mathbb{P}(\delta^{-}\leq\ell)\;\geq\;\xi^{as}$ (3)
$\;\;\mathbb{P}(\ell<f(X)<h)\;\leq\;\rho$ (4)
@@
We note that there is a trade-off between the rate of acceptance of unsafe devices (first term in the objective function) and the rate of acceptance of safe devices (2). Similarly, there is a trade-off between the rate of rejection of safe devices (second term in the objective function) and the rate of rejection of unsafe devices (3). The existence of these two constraints imposes minimum performance guarantees. The lack of these constraints, along with the absence of significant limitations on the FDA’s workload, results in a policy that defers almost all devices to minimize the objective function.
@@ 4.2 Structural Properties
In this section, we derive important structural properties of the optimization problem, allowing us to find a closed-form solution in many instances and develop an efficient heuristic to find near-optimal results in other instances.
@@
$\min_{\ell,h}$ $\;\;\lambda\;\mathbb{P}(\delta^{+}\leq\ell)+(1-\lambda)\;\mathbb{P}(\delta^{-}\geq h)$ (6)
s.t. $\;\;\mathbb{P}(\delta^{+}\geq h)\;\geq\;\xi^{ru}$ (7)
$\;\;\mathbb{P}(\delta^{-}\leq\ell)\;\geq\;\xi^{as}$ (8)
@@
To solve the Relaxed Problem, we introduce an auxiliary problem for which we can derive a closed-form solution. We then demonstrate how this auxiliary problem is in essence equivalent to the relaxed problem. We define the auxiliary problem as follows:
@@
where $\theta\in(0,1)$. Also, $h(\xi^{ru})\;=\;\sup\left\{h\in[0,1]:\;\mathbb{P}(\delta^{+}\geq h)\;\geq\;\xi^{ru}\right\}$ denotes the highest value of $h$ for which the true positive rate is greater than or equal to $\xi^{ru}$, and $\ell(\xi^{as})\;=\;\inf\left\{\ell\in[0,1]:\;\mathbb{P}(\delta^{-}\leq\ell)\;\geq\;\xi^{as}\right\}$ denotes the smallest value of $\ell$ for which the true negative rate is greater than or equal to $\xi^{as}$.
We next derive a closed-form solution for the Auxiliary Problem.
@@
    if $h(\xi^{ru})\;>\;\ell(\xi^{as})$, then the two-threshold optimal solution of the Auxiliary Problem is optimal in the Relaxed Problem,
@@
4.3 Nested Search
Given the structural properties derived in the prior section, finding a closed-form solution to the optimization problem in Section 4.1 remains challenging due to the workload constraint, which couples the computations of the low and high thresholds. In this section, we leverage the structural properties of our problem and develop a nested search approach to find thresholds $\ell$ and $h$.
We begin by reformulating the workload constraint (4). Let $N^{+}$ and $N^{-}$ denote the total number of unsafe and safe devices in the training data, respectively. Consequently, the total size of the training data is $N=N^{+}+N^{-}$, and the proportion of unsafe devices is $q=N^{+}/N$. Then, we have:
$\mathbb{P}(\ell<f(X)<h)$ $=1-\left(\mathbb{P}(f(X)<\ell)+\mathbb{P}(f(X)>h)\right),$
where,
$\mathbb{P}(f(X)<\ell)=q\;\mathbb{P}(\delta^{+}\leq\ell)+(1-q)\;\mathbb{P}(\delta^{-}\leq\ell),$
and,
$\mathbb{P}(f(X)>h)=q\;\mathbb{P}(\delta^{+}\geq h)+(1-q)\;\mathbb{P}(\delta^{-}\geq h).$
Let $\phi^{+}(\cdot)$ and $\phi^{-}(\cdot)$ be the empirical cumulative distributions of $\delta^{+}$ and $\delta^{-}$, respectively. Then, the workload constraint (4) can be written as
$q\left(\phi^{+}(h)-\phi^{+}(\ell)\right)+(1-q)\left(\phi^{-}(h)-\phi^{-}(\ell)\right)\;\leq\;\rho.$ (14)
We note that increasing $l$ weakly decreases the left-hand side of the workload constraint while weakly increasing the first term in the objective function. Similarly, decreasing $h$ weakly decreases the left-hand side of the workload constraint while weakly increasing the second term in the objective function. Thus, when $h(\xi^{ru})>\ell(\xi^{as})$ and $q\left(\phi^{+}(h(\xi^{ru}))-\phi^{+}(\ell(\xi^{as}))\right)+(1-q)\left(\phi^{-}(h(\xi^{ru}))-\phi^{-}(\ell(\xi^{as}))\right)\;\leq\;\rho$, the optimal thresholds should ensure that the workload constraint holds with equality. Under those conditions, the main optimization problem in Section 4.1 can be reformulated as follows:
$\min_{\ell,h}$ $\;\;\lambda\;\phi^{+}(\ell)+(1-\lambda)\;\left(1-\phi^{-}(h)\right)$
s.t. $\;\;q\left(\phi^{+}(h)-\phi^{+}(\ell)\right)+(1-q)\left(\phi^{-}(h)-\phi^{-}(\ell)\right)=\rho$
$\;\;\ell(\xi^{as})\;\leq\;\ell\;\leq\;h\;\leq\;h(\xi^{ru}).$
For the above optimization, it is still not possible to compute a closed-form solution as the workload constraint is coupled with the objective. However, we can use a nested search approach to identify the optimal thresholds $\ell$ and $h$. Specifically, we treat $\ell$ as the primary decision variable. For each candidate $\ell$, we then solve for a unique $h^{*}(\ell)$ that enforces the workload constraint in equality (provided such $h$ exists). The procedure proceeds in two key steps:
Step 1 (Inner sub-problem). For a given $\ell\in\bigl{[}\ell(\xi^{as}),\,h(\xi^{ru})\bigr{]}$, we seek $h^{*}(\ell)\in[\ell,\,h(\xi^{ru})]$ such that
$q\,\bigl{(}\phi^{+}\bigl{(}h^{*}(\ell)\bigr{)}-\phi^{+}(\ell)\bigr{)}\;+\;\bigl{(}1-q\bigr{)}\,\bigl{(}\phi^{-}\bigl{(}h^{*}(\ell)\bigr{)}-\phi^{-}(\ell)\bigr{)}\;=\;\rho\,,$ (15)
where for some large values of $\ell$, the above equation might be infeasible. To find such a solution, we employ a bisection search as follows:
  1. 1.
    Initialize $\bigl{[}h_{L},h_{R}\bigr{]}=\bigl{[}\ell,h(\xi^{ru})\bigr{]}$.
  2. 2.
    Define
    $g(\ell,h)\;=\;q\,\bigl{(}\phi^{+}(h)-\phi^{+}(\ell)\bigr{)}\;+\;\bigl{(}1-q\bigr{)}\,\bigl{(}\phi^{-}(h)-\phi^{-}(\ell)\bigr{)}-\rho\,.$
  3. 3.
    While $\,|h_{R}-h_{L}|$ is above a tolerance $\zeta$, evaluate $\,h_{\mathrm{mid}}=\tfrac{1}{2}(h_{L}+h_{R})\,$ and then:
    $\begin{cases}h_{L}=h_{\mathrm{mid}},&\text{if }g\bigl{(}\ell,h_{\mathrm{mid}}\bigr{)}<0\,,\\[3.0pt] h_{R}=h_{\mathrm{mid}},&\text{otherwise}.\end{cases}$
  4. 4.
    The final solution, $h^{*}_{\zeta}(\ell)$, is the midpoint of $\bigl{[}h_{L},h_{R}\bigr{]}$.
Step 2 (Outer sub-problem). We next identify the optimal $\ell$ within the interval $\bigl{[}\ell(\xi^{as}),\,h(\xi^{ru})\bigr{]}$. We discretize $\ell$ over a uniform grid:
=$\mathcal{L}\;=\;\{\ell_{1},\,\ell_{2},\dots,\ell_{m}\}\;\subseteq\;[\ell(\xi^{as}),\,h(\xi^{ru})),\quad\text{with }m\;=\;\left\lceil\frac{\,h(\xi^{ru})-\ell(\xi^{as})\,}{\Delta}\right\rceil.$
We proceed as follows for each $\ell\in\mathcal{L}$:
  1. 1.
    Apply the bisection method to compute $h^{*}_{\zeta}(\ell)$. If such a solution exists, add $\ell$ to the set of feasible candidates $\mathcal{L}_{\text{feas}}$; otherwise, discard $\ell$.
  2. 2.
    Evaluate the objective and record it:
    $F\bigl{(}\ell\bigr{)}\;=\;\lambda\;\phi^{+}(\ell)+(1-\lambda)\;\left(1-\phi^{-}(h^{*}_{\zeta}(\ell))\right).$
Finally, select the thresholds as
$\ell^{NS}\;=\;\arg\min_{\ell\in\mathcal{L}_{\text{feas}}}F\big{(}\ell\big{)},\quad h^{NS}\;=\;h^{*}_{\zeta}\big{(}\ell^{NS}\big{)},$
and return $\bigl{(}\ell^{NS},h^{NS}\bigr{)}$ as the solution. The overall procedure is summarized in Algorithm 1.
Algorithm 1 Nested Search Algorithm
1:Input: Empirical CDFs $\phi^{+}$ and $\phi^{-}$; parameters $q$, $\lambda$, $\rho$, $\xi^{as}$, $\xi^{ru}$; grid step $\Delta$; tolerance $\zeta$
2:Initialize: Set of feasible candidates $\mathcal{L}_{\text{feas}}\leftarrow\emptyset$
3:Construct a grid $\mathcal{L}$ over $[\ell(\xi^{as}),h(\xi^{ru}))$ using the grid step $\Delta$
4:for each $\ell\in\mathcal{L}$ do
5:     Set $h_{L}\leftarrow\ell$, $h_{R}\leftarrow h(\xi^{ru})$
6:     Compute $g(\ell,h_{L})$ and $g(\ell,h_{R})$, where $g(\ell,h)=q(\phi^{+}(h)-\phi^{+}(\ell))+(1-q)(\phi^{-}(h)-\phi^{-}(\ell))-\rho$
7:     if $g(\ell,h_{L})>0$ or $g(\ell,h_{R})<0$ then
8:         continue to next $\ell$
9:     while $|h_{R}-h_{L}|>\zeta$ do
10:         $h_{\text{mid}}\leftarrow(h_{L}+h_{R})/2$
11:         Compute $g(\ell,h_{\text{mid}})$
12:         if $g(\ell,h_{\text{mid}})<0$ then
13:              $h_{L}\leftarrow h_{\text{mid}}$
14:         else
15:              $h_{R}\leftarrow h_{\text{mid}}$
16:     Set $h^{*}_{\zeta}(\ell)\leftarrow(h_{L}+h_{R})/2$
17:     Add $\ell$ to the set of feasible candidates $\mathcal{L}_{\text{feas}}$
18:     Compute $F(\ell)=\lambda\phi^{+}(\ell)+(1-\lambda)(1-\phi^{-}(h^{*}_{\zeta}(\ell)))$ and record it
19:Select the thresholds as:
$\ell^{NS}=\arg\min_{\ell\in\mathcal{L}_{\text{feas}}}F\big{(}\ell\big{)},\quad h^{NS}=h^{*}_{\zeta}\big{(}\ell^{NS}\big{)}$
Utilizing the structural properties outlined in Section 4.2 and our nested search approach, we present our main theoretical result, which introduces a systematic approach to solving the optimization problem (1)-(5).
Theorem 4.3
For the optimization problem (1)-(5), we have: